OpenStudy (anonymous):

y= - 2 - x–3y+2z=9 3x+3y–3z=3

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

can someone help!

OpenStudy (alexandervonhumboldt2):

i dont undertsnad what they mean

OpenStudy (anonymous):

y=-2-x and so on

OpenStudy (alexandervonhumboldt2):

y=-2-x–3y+2z=9 3x+3y–3z=3 like this? and what do we need to do?

OpenStudy (anonymous):

yes we need to solve for x?

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (alexandervonhumboldt2):

y=-2-x–3y+2z=9 3x+3y–3z=3 substitue: 3x+27-3z=3

OpenStudy (anonymous):

okay

OpenStudy (alexandervonhumboldt2):

i got confused

OpenStudy (alexandervonhumboldt2):

i get no answer

OpenStudy (anonymous):

oooohh

OpenStudy (anonymous):

it's okay